OpenStudy (anonymous):

lim [x^3 + 8]/ [x^2 + 7x + 10] x->-2

OpenStudy (anonymous):

\[\lim_{x \rightarrow -2} \frac{(x+2)(x^2-2x+4)}{(x+5)(x+2)}=\lim_{x \rightarrow -2} \frac{x^2-2x+4}{x+5}\]

OpenStudy (anonymous):

plug in -2 now, you get your limit

OpenStudy (anonymous):

plug in -2 for each x.. u will get the answer!!